# Continuum Transformers Perform In-Context Learning by Operator Gradient Descent

### v1

@@ Abstract
Transformers robustly exhibit the ability to perform in-context learning, whereby their predictive accuracy on a task can increase not by parameter updates but merely with the placement of training samples in their context windows. Recent works have shown that transformers achieve this by implementing gradient descent in their forward passes. Such results, however, are restricted to standard transformer architectures, which handle finite-dimensional inputs. In the space of PDE surrogate modeling, a generalization of transformers to handle infinite-dimensional function inputs, known as “continuum transformers,” has been proposed and similarly observed to exhibit in-context learning. Despite impressive empirical performance, such in-context learning has yet to be theoretically characterized. We herein demonstrate that continuum transformers perform in-context operator learning by performing gradient descent in an operator RKHS. We demonstrate this using novel proof strategies that leverage a generalized representer theorem for Hilbert spaces and gradient flows over the space of functionals of a Hilbert space. We additionally show the operator learned in context is the Bayes Optimal Predictor in the infinite depth limit of the transformer. We then provide empirical validations of this optimality result and demonstrate that the parameters under which such gradient descent is performed are recovered through the continuum transformer training.
@@ 1 Introduction
While LLMs were the first setting in which in-context learning abilities of transformers were exploited, interest is increasing in its broader application to other sequence prediction tasks. In particular, a seemingly orthogonal subfield of machine learning is that of accelerating the solution of partial differential equations (PDEs) [6, 7, 8, 9, 10, 11]. In this setting, the sequence is no longer of finite-dimensional vectors, but instead of infinite-dimensional functions. Nonetheless, transformers have been adapted to this setting, with an architecture known as “continuum transformers” [12]. More surprising still is that such transformers continue to exhibit in-context learning, whereby related PDEs can be efficiently solved with the similar placement of solution pairs in the continuum transformer context window [13, 14, 15, 16].
Unlike in the finite-dimensional setting, this generalized, functional in-context learning yet remains to be theoretically characterized. We, therefore, herein extend this line of theoretical inquiry to characterize the continuum transformers that have been leveraged for in-context operator learning [12]. In particular, our contributions are as follows:
  • Proving continuum transformers perform in-context operator learning by performing gradient descent in a Reproducing Kernel Hilbert Space over operators and that the resulting in-context predictor recovers the Bayes Optimal Predictor under well-specified parameter choices of the transformer. Such proofs make novel use of a generalized representer theorem for Hilbert spaces and Gaussian measures over Hilbert spaces.
@@ 2.1 Neural Operators
Neural operator methods, while more broadly applicable, most often seek to amortize the solution of a spatiotemporal PDE. Such PDEs are formally described by spatial fields that evolve over time, namely some $u:\Omega\times[0,\infty)\rightarrow\mathbb{R}$, where $x\in\Omega\subset\mathbb{R}^{d}$ are spatial coordinates and $t\in[0,\infty)$ is a time coordinate. While the more abstract operator learning framework can be formulated as seeking to learn a map $\widehat{\mathcal{G}}:\mathcal{A}\rightarrow\mathcal{U}$ between two function spaces $\mathcal{A}$ and $\mathcal{U}$, we are most often interested in learning time-rollout maps, in which the input and output function spaces are identical. In such cases, it is assumed a dataset of the form $\mathcal{D}:=\{(u^{(0)}_{i},u^{(T)}_{i})\}$ is available, where $u^{(0)}:\Omega\rightarrow\mathbb{R}$ is the initial condition, for which there exists some true operator $\mathcal{G}$ such that, for all $i$, $u_{i}^{(T)}=\mathcal{G}(u_{i}^{(0)})$.
While many different learning-based approaches have been proposed to solve this operator learning problem, they all can be abstractly framed as seeking
@@
Notably, in such works, it is often assumed that the operator of interest is mapping between two sufficiently smooth spaces, on which the PDE is well-defined [17, 18]. That is, most often $\mathcal{A}=\mathcal{U}=\mathcal{H}^{s}(\Omega)$, where $\mathcal{H}^{s}(\Omega):=\{f:||f||_{\mathcal{H}^{s}(\Omega)}<\infty\}$ is the standard $s$-Sobolev space, in which $||f||_{\mathcal{H}^{s}(\Omega)}:=\sum_{|\alpha|\leq k:\alpha\in\mathbb{N}^{d}}||D^{\alpha}f||_{\mathcal{L}^{2}(\Omega)}$. $s$ is the smoothness parameter as determined by the PDE, and $\alpha$ is the condensed notation for the corresponding mixed partials, i.e. $\alpha=(\alpha_{1},...,\alpha_{d})$ corresponds to $D^{\alpha}:=\partial^{\alpha_{1}}_{x_{1}}...\partial^{\alpha_{d}}_{x_{d}}$.
Within this broad family of operator learning, the most widely employed classes are the Deep Operator Networks (DeepONets) [19, 20, 21, 22] and Fourier Neural Operators (FNOs) [23, 24, 6, 25]. FNOs are parameterized as a sequence of layers of linear operators, given as kernel integral transforms, with standard intermediate ReLU nonlinearities. Formally, a single layer is then given by
[…]
where the latter equivalence follows from the standard convolution theorem under the assumption the learned kernel is translation invariant, i.e. $k_{\ell}(x,y)=k^{\prime}_{\ell}(x-y)$, where $\mathcal{F}$ denotes the Fourier transform. Due to improved computational efficiency, therefore, the learning of the kernel is done in Fourier rather than real space, where $R_{\ell}$ is learned up to some fixed truncation point $k_{\max}$.
@@ 2.2 In-Context Learning and Continuum Transformers
The standard attention mechanism is parameterized by $\theta:=\{W_{k},W_{q},W_{v}\}$, where $W_{k}\in\mathbb{R}^{d_{k}\times d}$, $W_{q}\in\mathbb{R}^{d_{q}\times d}$, and $W_{v}\in\mathbb{R}^{d_{v}\times d}$ for sequential data $X\in\mathbb{R}^{d\times n}$ [26]. In a generalized form,
@@
where $M\in\mathbb{R}^{n\times n}$ is a masking matrix and $H:\mathbb{R}^{d_{q}\times n}\times\mathbb{R}^{d_{k}\times n}\rightarrow\mathbb{R}^{n\times n}$ is a nonlinear transform of key-query similarity measures. This resulting matrix $H(Q,K)$ is referred to as the “attention weights” matrix. Most often, $d_{k}=d_{q}$ and $H:=\mathrm{softmax}$, i.e. $[H(Q,K)]_{i,j}=\exp(Q^{\top}_{i}K_{j})/\sum_{\ell}\exp(Q^{\top}_{i}K_{\ell})$ [27]. Transformers are repeated compositions of such attention blocks with residual connections and feedforward and normalization layers.
We now highlight the “continuum attention” mechanism proposed in [12] that generalizes transformers to operator learning; “continuum” here highlights that such an architecture models data in its discretization-agnostic function form as with other neural operator methods. In particular, in place of $x_{i}\in\mathbb{R}^{d}$, $x_{i}\in\mathcal{X}$ for some Hilbert space $\mathcal{X}$; notably, the presentation of [12] was particular to $\mathcal{X}=\mathcal{F}(\Omega;\mathbb{R}^{d})$, i.e. the set of all functions $x:\Omega\rightarrow\mathbb{R}^{d}$. The natural generalization of the attention mechanism for function inputs then replaces the $W_{k},W_{q}$, and $W_{v}$ matrices with linear operators. For a sequence $x\in\mathcal{X}^{n}$, the key, query, and value operators respectively map $\mathcal{W}_{k}:\mathcal{X}^{n}\rightarrow\mathcal{K}^{n}$, $\mathcal{W}_{q}:\mathcal{X}^{n}\rightarrow\mathcal{Q}^{n}$, and $\mathcal{W}_{v}:\mathcal{X}^{n}\rightarrow\mathcal{V}^{n}$, where $\mathcal{K},\mathcal{Q},$ and $\mathcal{V}$ are Hilbert spaces. In practice, such operators are implemented as kernel integral transforms, as $\mathcal{W}_{q}x_{i}=\mathcal{F}^{-1}(R_{q}\odot\mathcal{F}x_{i}),$ where $R_{q}$ is the Fourier parameterization of the query kernel, with $R_{k}$ and $R_{v}$ similarly defined for the key and value kernels. The continuum attention mechanism assumes $\mathcal{K}=\mathcal{Q}$ and defines
@@
We then suppose $y_{i}^{(t)}=f_{i}(x_{i}^{(t)})$, where $f_{i}\neq f_{j}$ for $i\neq j$. This is the critical difference in the in-context learning setting versus typical learning settings: in the standard setting, $f_{i}=f$ is fixed across samples and necessarily matches the $f^{\prime}$ present at test time, meaning the goal for the learner is to merely learn such an $f$. In in-context learning, however, the learning algorithm must be capable of “learning” an unseen $f^{\prime}$ at inference time without parameter updates.
@@ 2.3 Operator Meta-Learning
Significant interest has emerged in meta-operator learning for spatiotemporal PDEs, in which a single network maps from initial conditions to final states across system specifications [28, 29, 30, 31, 13, 32, 33, 14]. Formally, unlike the traditional setting discussed in Section 2.1, here the dataset consists of sub-datasets $\mathcal{D}:=\cup_{i}\mathcal{D}_{i}$, where $\mathcal{D}_{i}:=\{(u_{i;0}^{(j)},u_{i;T}^{(j)})\}^{n_{i}}_{j=1}$, for which the true operator can vary across sub-datasets, i.e. $\mathcal{G}_{i}$ satisfies $u_{i;T}^{(j)}=\mathcal{G}_{i}(u_{i;0}^{(j)})$ $\forall j\in[n_{i}]$ but $\mathcal{G}_{i}\neq\mathcal{G}_{i^{\prime}}$ for $i\neq i^{\prime}$. The goal is to then, given only a limited number of training samples $\widetilde{\mathcal{D}}:=\{(\widetilde{u}_{0}^{(j)},\widetilde{u}_{T}^{(j)})\}^{\widetilde{n}}_{j=1}$ with $\widetilde{n}\ll\sum_{i}n_{i}$ for a fixed, potentially unseen operator $\mathcal{G}^{\prime}$, learn an approximation $\widehat{\mathcal{G}}\approx\widetilde{\mathcal{G}}$. Most often, this is done by pre-training a meta-learner $\mathcal{G}_{\mathrm{ML}}$ on $\mathcal{D}$ and then fine-tuning $\mathcal{G}_{\mathrm{ML}}$ on $\widetilde{\mathcal{D}}$ to arrive at $\widetilde{\mathcal{G}}$.
Building off of the observed ICL of transformers, the fine-tuning of these meta-learners too can be performed via explicit weight modification or via in-context learning. In many operator learning settings, in-context learning is the dominant paradigm for fine-tuning. In fact, an entire offshoot of meta-learning known as in-context operator networks (ICONs) has spawned from this approach [13, 14, 15, 16]. Recent works in this vein have demonstrated notable empirical performance leveraging bespoke transformer architectures built atop the continuum attention from Equation 4 [13, 34, 35, 36]. Such behavior has been empirically observed but has yet to be theoretically characterized.
2.4 Related Works
We are interested herein in providing a theoretical characterization of the in-context learning exhibited by continuum attention-based ICONs, paralleling that done for finite-dimensional transformers [3, 4, 5]. Most relevant in this line of work is that of [27]. Loosely speaking, they demonstrated that, if a kernel $\kappa(\cdot,\cdot)$ is defined with $\mathbb{H}$ denoting its associated RKHS and a transformer is then defined with a specific $W_{k},W_{q},W_{v}$ and $[H(Q,K)]_{i,j}=\kappa(Q_{i},K_{j})$ from Equation 3, an inference pass through $m$ layers of such a transformer predicts $y^{(n+1)}_{i}$ from Equation 5 equivalently to $\widehat{f}(x^{(n+1)}_{i})$, where $\widehat{f}$ is the model resulting from $m$ steps of functional gradient descent in $\mathbb{H}$. They additionally demonstrated that the predictor learned by such ICL gradient descent is Bayes optimal under particular circumstances. These results are formally provided in Appendix A. Notably, these results require non-trivial changes to be generalized to operator RKHSs as we do herein. While previous works have yet to formally characterize ICL for continuum transformers, a recent work [37] began a line of inquiry in this direction. Their work, however, studied a fundamentally distinct aspect of functional ICL than that studied herein, characterizing its sample complexity and studying its resulting generalization, specifically in the restricted setting of linear elliptic PDEs.
@@ 3 Theoretical Characterization
3.1 In-Context Learning Setup
We now consider a generalization of the continuum attention, paralleling that of Equation 3, allowing for more general key-query similarity measures. In particular, instead of restricting $\mathcal{Q}=\mathcal{K}$ as required for Equation 4 to be well-defined, we allow $H:\mathcal{Q}^{n+1}\times\mathcal{K}^{n+1}\to(\mathcal{L}(\mathcal{V}))^{(n+1)\times(n+1)}$, where $\mathcal{L}(\mathcal{V})$ denotes the set of bounded linear operators from $\mathcal{V}$ to $\mathcal{V}$; note that the “$n+1$” convention is adopted as data is of the form given by Equation 5. This generalization subsumes the softmax form assumed in Equation 4, where for $c\in\mathbb{R}$, $c\in\mathcal{L}(V)$ is understood to be defined as $cf_{v}$ for $f_{v}\in\mathcal{V}$, from which we can similarly view $\mathrm{softmax}:\mathcal{Q}^{n+1}\times\mathcal{K}^{n+1}\to(\mathcal{L}(\mathcal{V}))^{(n+1)\times(n+1)}$. If $\mathcal{V}=\mathcal{X}$, an $m$-layer continuum transformer $\mathcal{T}:\mathcal{X}^{n+1}\rightarrow\mathcal{X}^{n+1}$ consisting of generalized continuum attention layers with residual connections is well-defined as $\mathcal{T}:=\mathcal{T}_{m}\circ...\circ\mathcal{T}_{0}$, where

### v2

@@ Abstract
Transformers robustly exhibit the ability to perform in-context learning, whereby their predictive accuracy on a task can increase not by parameter updates but merely with the placement of training samples in their context windows. Recent works have shown that transformers achieve this by implementing gradient descent in their forward passes. Such results, however, are restricted to standard transformer architectures, which handle finite-dimensional inputs. In the space of PDE surrogate modeling, a generalization of transformers to handle infinite-dimensional function inputs, known as “continuum transformers,” has been proposed and similarly observed to exhibit in-context learning. Despite impressive empirical performance, such in-context learning has yet to be theoretically characterized. We herein demonstrate that continuum transformers perform in-context operator learning by performing gradient descent in an operator RKHS. We demonstrate this using novel proof strategies that leverage a generalized representer theorem for Hilbert spaces and gradient flows over the space of functionals on a Hilbert space. We further show the operator learned in context is the Bayes Optimal Predictor in the infinite depth limit of the transformer. We then provide empirical validations of this result and demonstrate that the parameters under which such gradient descent is performed are recovered through pre-training.
@@ 1 Introduction
While LLMs were the first setting in which in-context learning abilities of transformers were exploited, interest is increasing in its broader application to other sequence prediction tasks. In particular, a seemingly orthogonal subfield of machine learning is that of accelerating the solution of partial differential equations (PDEs) [6, 7, 8, 9, 10, 11]. In this setting, the sequence is no longer of finite-dimensional vectors, but instead of infinite-dimensional functions. Nonetheless, transformers have been adapted to this setting, with an architecture known as “continuum transformers” [12]. More surprising still is that such transformers continue to exhibit in-context learning, whereby related PDEs can be efficiently solved with the similar placement of solution pairs in the continuum transformer context window [13, 14, 15].
Unlike in the finite-dimensional setting, this generalized, functional in-context learning yet remains to be theoretically characterized. We, therefore, herein extend this line of theoretical inquiry to characterize the continuum transformers that have been leveraged for in-context operator learning [12]. Our contributions, therefore, are twofold: the insights afforded by such theoretical analysis and the development of a mathematical framework for doing such analysis, as we highlight in greater detail in the main text. In particular, our contributions are as follows:
  • Proving continuum transformers perform in-context operator learning by performing gradient descent in a Reproducing Kernel Hilbert Space of operators and that the resulting in-context predictor recovers the Bayes Optimal Predictor under well-specified parameter choices of the transformer. Such proofs required the modeling of a generalized continuum attention mechanism and subsequent novel usages of a generalized representer theorem for Hilbert spaces and Gaussian measures over Hilbert spaces.
@@ 2.1 Neural Operators
Neural operator methods, while more broadly applicable, most often seek to amortize the solution of a spatiotemporal PDE. Such PDEs are formally described by spatial fields that evolve over time, namely some $u:\Omega\times[0,\infty)\rightarrow\mathbb{R}$, where $x\in\Omega\subset\mathbb{R}^{d}$ are spatial coordinates and $t\in[0,\infty)$ is a time coordinate. While the more abstract operator learning framework can be formulated as seeking to learn a map $\widehat{\mathcal{G}}:\mathcal{A}\rightarrow\mathcal{U}$ between two function spaces $\mathcal{A}$ and $\mathcal{U}$, we are most often interested in learning time-rollout maps, in which the input and output function spaces are identical. In such cases, it is assumed a dataset of the form $\mathcal{D}:=\{(u^{(0)}_{i},u^{(T)}_{i})\}$ is available, where $u^{(0)}:\Omega\rightarrow\mathbb{R}$ is the initial condition, for which there exists some true operator $\mathcal{G}$ such that, for all $i$, $u_{i}^{(T)}=\mathcal{G}(u_{i}^{(0)})$. While many different learning-based approaches have been proposed to solve this operator learning problem, they all can be abstractly framed as seeking
@@
Within this broad family of operator learning, the most widely employed classes are the Deep Operator Networks (DeepONets) [16, 17, 18, 19] and Fourier Neural Operators (FNOs) [20, 21, 6, 22]. FNOs are parameterized as a sequence of layers of linear operators, given as kernel integral transforms, with standard intermediate ReLU nonlinearities. Formally, a single layer is then given by $\sigma(\mathcal{F}^{-1}(R_{\ell}\odot\mathcal{F}(u)))$, where $\mathcal{F}$ denotes the Fourier transform and $R_{\ell}$ the learnable kernel parameter. This layer encodes a translationally invariant kernel integral transform.
@@ 2.2 In-Context Learning and Continuum Transformers
The standard attention mechanism is parameterized by $\theta:=\{W_{k},W_{q},W_{v}\}$, where $W_{k}\in\mathbb{R}^{d_{k}\times d}$, $W_{q}\in\mathbb{R}^{d_{q}\times d}$, and $W_{v}\in\mathbb{R}^{d_{v}\times d}$ for sequential data $X\in\mathbb{R}^{d\times n}$ [23]. Formally,
@@
where $M\in\mathbb{R}^{n\times n}$ is a masking matrix and $H:\mathbb{R}^{d_{q}\times n}\times\mathbb{R}^{d_{k}\times n}\rightarrow\mathbb{R}^{n\times n}$ is a nonlinear transform of key-query similarity measures. This resulting matrix $H(Q,K)$ is referred to as the “attention weights” matrix. Most often, $d_{k}=d_{q}$ and $H:=\mathrm{softmax}$, i.e. $H(Q,K)]_{i,j}=\exp(Q^{\top}_{i}K_{j})/\sum_{\ell}\exp(Q^{\top}_{i}K_{\ell}) [$24]. Transformers are repeated compositions of such attention blocks with residual connections and feedforward and normalization layers.
We now highlight the “continuum attention” mechanism proposed in [12] that generalizes transformers to operator learning; “continuum” here highlights that such an architecture models data in its discretization-agnostic function form as with other neural operator methods. In particular, in place of $x_{i}\in\mathbb{R}^{d}$, $x_{i}\in\mathcal{X}$ for some Hilbert space $\mathcal{X}$. The natural generalization of the attention mechanism for function inputs then replaces the $W_{k},W_{q}$, and $W_{v}$ matrices with linear operators. For a sequence $x\in\mathcal{X}^{n}$, the key, query, and value operators respectively map $\mathcal{W}_{k}:\mathcal{X}^{n}\rightarrow\mathcal{K}^{n}$, $\mathcal{W}_{q}:\mathcal{X}^{n}\rightarrow\mathcal{Q}^{n}$, and $\mathcal{W}_{v}:\mathcal{X}^{n}\rightarrow\mathcal{V}^{n}$, where $\mathcal{K},\mathcal{Q},$ and $\mathcal{V}$ are Hilbert spaces. Notably, the original framing of continuum transformers [12] generalizes attention by considering an infinite index set, rather than considering infinite-dimensional tokens as we do here; we discuss how these two seemingly distinct notions are equivalent in Appendix A. In practice, such operators are implemented as kernel integral transforms, as $\mathcal{W}_{q}x_{i}=\mathcal{F}^{-1}(R_{q}\odot\mathcal{F}x_{i}),$ where $R_{q}$ is the Fourier parameterization of the query kernel, with $R_{k}$ and $R_{v}$ similarly defined for the key and value kernels. The continuum attention mechanism assumes $\mathcal{K}=\mathcal{Q}$ and defines
@@
We then suppose $y_{i}^{(t)}=f_{i}(x_{i}^{(t)})$, where $f_{i}\neq f_{j}$ for $i\neq j$. This is the critical difference in the in-context learning setting: in the standard setting, $f_{i}=f$ is fixed across samples and matches the $f^{\prime}$ at test time, meaning the goal for the learner is to learn such an $f$. In ICL, however, the learning algorithm must be capable of “learning” an unseen $f^{\prime}$ at inference time without parameter updates.
@@ 2.3 Operator Meta-Learning
Significant interest has emerged in meta-operator learning for spatiotemporal PDEs, in which a single network maps from initial conditions to final states across system specifications [25, 26, 27, 28, 13, 29]. Formally, unlike the traditional setting discussed in Section 2.1, here the dataset consists of sub-datasets $\mathcal{D}:=\cup_{i}\mathcal{D}_{i}$, where $\mathcal{D}_{i}:=\{(u_{i;0}^{(j)},u_{i;T}^{(j)})\}^{n_{i}}_{j=1}$, for which the true operator can vary across sub-datasets, i.e. $\mathcal{G}_{i}$ satisfies $u_{i;T}^{(j)}=\mathcal{G}_{i}(u_{i;0}^{(j)})$ $\forall j\in[n_{i}]$ but $\mathcal{G}_{i}\neq\mathcal{G}_{i^{\prime}}$ for $i\neq i^{\prime}$. The goal is to then, given only a limited number of training samples $\widetilde{\mathcal{D}}:=\{(\widetilde{u}_{0}^{(j)},\widetilde{u}_{T}^{(j)})\}^{\widetilde{n}}_{j=1}$ with $\widetilde{n}\ll\sum_{i}n_{i}$ for a fixed, potentially unseen operator $\mathcal{G}^{\prime}$, learn an approximation $\widehat{\mathcal{G}}\approx\widetilde{\mathcal{G}}$. Most often, this is done by pre-training a meta-learner $\mathcal{G}_{\mathrm{ML}}$ on $\mathcal{D}$ and then fine-tuning $\mathcal{G}_{\mathrm{ML}}$ on $\widetilde{\mathcal{D}}$ to arrive at $\widetilde{\mathcal{G}}$.
Building off of the observed ICL of transformers, the fine-tuning of these meta-learners too can be performed via explicit weight modification or via in-context learning. An entire offshoot of meta-learning known as in-context operator networks (ICONs) has spawned from the latter [13, 14, 15, 30]. Recent works in this vein have demonstrated notable empirical performance leveraging bespoke transformer architectures built atop the continuum attention from Equation 3 [13, 31, 32, 33]. Such behavior has been observed but has yet to be theoretically characterized.
@@ 3 Theoretical Characterization
3.1 In-Context Learning Model
For quick reference, we provide a compilation of the notation used throughout this and the remaining sections in Appendix K. We now consider a generalization of the continuum attention, paralleling that of Equation 2, allowing for more general key-query similarity measures. In particular, instead of restricting $\mathcal{Q}=\mathcal{K}$ as required for Equation 3 to be well-defined, we allow $H:\mathcal{Q}^{n+1}\times\mathcal{K}^{n+1}\to(\mathcal{L}(\mathcal{V}))^{(n+1)\times(n+1)}$, where $\mathcal{L}(\mathcal{V})$ denotes the set of bounded linear operators from $\mathcal{V}$ to $\mathcal{V}$; note that the “n+1n+1” convention is adopted as data is of the form given by Equation 4. This generalization subsumes the softmax form assumed in Equation 3, where for $c\in\mathbb{R}$, $c\in\mathcal{L}(V)$ is understood to be defined as $cf_{v}$ for $f_{v}\in\mathcal{V}$, from which we can similarly view $\mathrm{softmax}:\mathcal{Q}^{n+1}\times\mathcal{K}^{n+1}\to(\mathcal{L}(\mathcal{V}))^{(n+1)\times(n+1)}$. If $\mathcal{V}=\mathcal{X}$, an $m$-layer continuum transformer $\mathcal{T}:\mathcal{X}^{n+1}\rightarrow\mathcal{X}^{n+1}$ consisting of generalized continuum attention layers with residual connections is well-defined as $\mathcal{T}:=\mathcal{T}_{m}\circ...\circ\mathcal{T}_{0}$, where